# ICONS: Influence Consensus for Vision-Language Data Selection

### v3

@@ Abstract
Training vision-language models via instruction tuning often relies on large mixtures of data spanning diverse tasks and domains. However, these mixtures frequently include redundant information, increasing computational costs without proportional performance gains—necessitating more effective data selection strategies. Existing methods typically rely on task-agnostic heuristics to estimate data importance or focus on optimizing single tasks in isolation, limiting their effectiveness in multitask settings. In this work, we introduce ICONS, a gradient-based Influence CONsensus approach for vision-language data Selection. Our method leverages first-order training dynamics to estimate the influence of individual training examples on validation performance and aggregates these estimates across tasks via majority voting over task-specific influences. This cross-task consensus identifies data points that are consistently valuable across tasks, enabling us to prioritize examples that drive overall performance. The voting-based design further mitigates issues such as score calibration and outlier sensitivity, resulting in robust and scalable data selection for diverse multitask mixtures. With only 20% of the data from LLaVA-665K and Cambrian-7M, our selected subsets retain 98.6% and 98.8% of the performance achieved with full datasets—and can even surpass full-data training at a 60% selection ratio on LLaVA-665K. Our approach also generalizes to unseen tasks and architectures, demonstrating strong transfer. We release two compact, high-utility subsets—LLaVA-ICONS-133K and Cambrian-ICONS-1.4M —preserving impactful training examples for efficient and scalable vision-language model development.
@@ 1 Introduction
Visual instruction tuning is a crucial step in training multimodal language models liu2024improved ; llava , enabling them to follow language instructions grounded in visual content. Recent approaches rely on large-scale datasets such as LLaVA-665K liu2024improved and Cambrian-7M tong2024cambrian , which contain 665K and 7M examples, respectively. While effective, these datasets introduce significant barriers to iteration and deployment: prolonged training times brown2020language ; kaplan2020scaling , high storage demands rajbhandari2020zero ; gadre2024datacomp , and substantial compute costs maini2024rephrasing ; wu2022sustainable . Moreover, not all examples contribute equally to all tasks—naively scaling up diverse data mixtures can introduce redundancy and inefficiency. This raises a fundamental question:
{mdframed}
[leftmargin=0pt, rightmargin=0pt, innerleftmargin=10pt, innerrightmargin=10pt, skipbelow=0pt]
Can we identify a compact, multitask-effective subset of training data that preserves model capabilities arcoss tasks while enabling faster experimentation?
Prior work has explored various data selection strategies, including gradient-based approaches xia2024less ; deng2024influential , influence functions yang2022dataset ; koh2017understanding , and diversity-based sampling yu2024diversify ; bukharin2023data . However, many of these methods either optimize for single tasks in isolation or maximize source diversity without aligning to downstream needs. In multitask visual instruction tuning, this is particularly limiting: optimizing for one task may hurt generalization, and task-agnostic diversity may dilute impact. Rather than selecting
@@
Figure 1: Influence consensus for vision-language data selection. (Left) Given a large scale visual instruction tuning dataset (LLaVA-665K), our method uses majority voting across task-specific influence scores to identify training samples that are consistently influential across multiple tasks, forming a compact 20% subset (LLaVA-ICONS-133K) with data points achieving influence consensus. (Right) The radar plot compares performance between LLaVA-665K and our selected subset, showing the selected subset achieves comparable results to the full dataset.
data based on per-task influence, we aim to identify samples that are broadly useful—training examples that consistently contribute across tasks. To do this, we aggregate gradient-based influence scores using a simple yet effective majority voting scheme.
@@
We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon the gradient-based selection approach LESS xia2024less . Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) uses influence consensus through majority voting to identify training samples that show consistent positive value across multiple tasks. This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum. While the computational cost of influence estimation is expensive, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate development of multimodal models, and enables reusable gradient datastores that amortize costs across iterations and deliver long-term savings.
@@
Using ICONS, we create LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, automatically curated 20% subsets of the LLaVA-665K dataset liu2024improved and Cambrian-7M tong2024cambrian dataset, respectively. These compact datasets maintain 98.6% and 98.8% of their original performance across multiple vision-language tasks, providing significant improvements over randomly selecting same-sized subsets (95.8% and 95.4%) and eliminating approximately two-thirds of the performance drop from shrinking the training data. Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original datasets at a 60% selection ratio for LLaVA-665K. Importantly, the selected subset shows strong transferability, e.g., LLaVA-ICONS-133K maintains 95.5-113.9% relative performance across unseen tasks, suggesting that ICONS identifies fundamentally valuable training data. We summarize our key contributions:
@@
    Our consensus-based selection outperforms all baselines (§3.2) and we ablate influence aggregation strategies and show the advantage of voting-based consensus (§3.3). We further show that ICONS exceeds 102% of full-dataset performance at a 60% selection ratio on LLaVA-665K3.5).
@@
    We release LLaVA-ICONS-133K and Cambrian-ICONS-1.4M, compact 20% subsets of LLaVA-665K and Cambrian-7M respectively, achieving near-full performance (98.6% and 98.8%), transferring well to unseen tasks (§3.4), and serving as standardized training sets for resource-efficient development.
2 Influence consensus for vision-language data selection
@@
Figure 2: ICONS. The Specialist stage (left) processes each task individually through three steps: (1) warmup training on a small subset of data, (2) gradient computation for both training and target task validation data, and (3) influence matrix computation to generate per-task influence scores. This process is repeated for each target task. The Generalist stage (right) performs Influence Consensus to aggregate information across tasks, where samples scoring above the $80^{\text{th}}$ percentile threshold for each task receive a vote. The final selection is made by summing votes across tasks and selecting the top 20% most influential samples, creating a compact yet highly effective training dataset that performs well across all tasks.
@@ 2.1 Problem formulation
Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})$ includes an image ${\bm{I}}_{i}$, natural language instruction ${\bm{x}}_{i}$, and corresponding target response ${\bm{y}}_{i}$ [^1] The framework supports multi-turn conversational data, yet we formalize the problem setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}),\quad\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})},$ (1)
where $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively. $\text{Score}(f,T_{k})$ is the task-specific evaluation score achieved by model $f$ on task $T_{k}$. We define the average relative performance across all tasks as Rel. = $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$. Rel. quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training coincide . Our objective is to select a subset where Rel. $\approx 1$, i.e., the model trained on the subset achieves comparable performance to the one trained with full dataset.
@@ 2.2 Preliminaries
Building on the problem formulation in §2.1, we formalize how to estimate the influence of training samples on downstream task performance. Since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn. 1, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator. Denote a training data point as ${\bm{z}}$ and a validation data point as ${\bm{z}}^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$. Following pruthi2020estimating ; xia2024less , we estimate how ${\bm{z}}\in\mathcal{D}$ affects validation loss by measuring its gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation. When training with SGD and batch size 1, using data point ${\bm{z}}$ at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate. To reduce the computational cost, we use the first-order Taylor expansion to estimate the loss on a given validation data point ${\bm{z}}^{\prime}$ at time step $t+1$ as:
@@
The influence of a training data point ${\bm{z}}$ on a validation data point ${\bm{z}}^{\prime}$ is:
$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
which we refer to as an influence score.
The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in xia2024less . through a greedy, first-order approximation, which leads to larger reductions in validation loss for point ${\bm{z}}^{\prime}$. While it omits second-order terms compared to influence functions koh2017understanding , it provides an efficient approximation to rank the impact of training samples xia2024less .
@@ 2.3 Specialist: individual task influence ranking
To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution. Following LESS xia2024less , the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
Step 1: Warm-up Training.  Following LESS xia2024less , we first apply LoRA hu2021lora on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5%) to obtain $f{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$. This allows the model to develop basic visual instruction-following capabilities.
@@
where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively. In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=Rg^{\prime}_{j}$, where $R\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability (johnson1984extensions, ). We further normalize the projected gradients, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences xia2024less .
Step 3: Influence matrix computation.  We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of the training data ${\bm{z}}_{i}$ on the validation data ${\bm{z}}^{\prime}_{j}$, and then the average influence of training data ${\bm{z}}_{i}$ on the target task $k$ is calculated as:
@@
This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence $\mathcal{T}_{k}$.
We can select a small training subset $\mathcal{S}_{k}$ for a given task $k$ by selecting the training examples ${\bm{z}}_{i}$ with the highest-influence values $\bar{I}_{k}({\bm{z}}_{i})$. This simple greedy approach has been shown by LESS to be successful, and thus we use it as our task-specific (“specialist”) baseline. However, recall that our goal is to select a single compact subset that maximizes the performance across all tasks. We address this disconnection between task-specific rankings and overall optimization by proposing a voting-based generalist approach to identify the most broadly impactful training data.

### v4

@@ Abstract
Training vision-language models via instruction tuning relies on large data mixtures spanning diverse tasks and domains, yet these mixtures frequently include redundant information that increases computational costs without proportional gains. Existing methods typically rely on task-agnostic heuristics to estimate data importance, limiting their effectiveness across tasks. We introduce ICONS, a gradient-based Influence CONsensus approach for vision-language data Selection. Our method leverages first-order training dynamics to estimate each example’s influence on validation performance, then aggregates these estimates across tasks via majority voting. This cross-task consensus identifies consistently valuable data points while mitigating score calibration and outlier sensitivity, enabling robust and scalable data selection for diverse multitask mixtures. Models trained on our selected 20% data subset from LLaVA-665K (respectively: from Cambrian-7M, from Vision-Flan-186K) retain 98.6% (respectively: 98.8%, 99.8%) of full-dataset performance. We demonstrate that our selected data generalizes to unseen tasks and model architectures, and release three compact subsets LLaVA-ICONS-133K, Cambrian-ICONS-1.4M, and Vision-Flan-ICONS-37K for efficient vision-language model development.
@@ 1 Introduction
Figure 1: Influence consensus for vision-language data selection. (Left) Given a large-scale visual instruction tuning dataset (LLaVA-665K), our method uses majority voting across task-specific influence scores to identify training samples that are consistently influential across multiple tasks, forming a compact 20% subset (LLaVA-ICONS-133K) with data points achieving influence consensus. (Right) The radar plot compares performance between LLaVA-665K and our selected subset, showing the selected subset achieves comparable results to the full dataset.
Visual instruction tuning (VIT) is a crucial step in training vision-language models (VLMs) [liu2024improved, llava], enabling them to follow language instructions grounded in visual content. Recent approaches rely on large-scale instruction following datasets such as LLaVA-665K [liu2024improved] and Cambrian-7M [tong2024cambrian]. While effective, these datasets introduce significant barriers to model training and deployment: prolonged training times [brown2020language, kaplan2020scaling], high storage demands [rajbhandari2020zero, gadre2024datacomp], and substantial compute costs [maini2024rephrasing, wu2022sustainable]. Moreover, not all training samples contribute equally to diverse downstream tasks and naively scaling up diverse data mixtures can introduce redundancy and inefficiency. This raises a fundamental question:
Can we identify a compact, multitask-effective subset of training data that preserves model capabilities across tasks while enabling faster experimentation?
Prior work has explored various data selection strategies, including gradient-based approaches [xia2024less, deng2024influential], influence functions [yang2022dataset, koh2017understanding], and diversity-based sampling [yu2024diversify, bukharin2023data]. However, many of these methods either optimize for single tasks in isolation or maximize source diversity without aligning to downstream needs. In the context of VIT data that supports diverse vision-language tasks, this is particularly limiting: optimizing for one task may hurt generalization, and task-agnostic diversity may dilute impact. Rather than selecting data based on per-task influence, we aim to identify samples that are broadly useful, with training examples that consistently contribute across tasks. To do this, we aggregate gradient-based influence scores using a simple yet effective majority voting approach to reach influence consensus.
@@
We introduce ICONS (Influence CONsensus vision-language data Selection), a method that builds upon LESS [xia2024less], a gradient-based selection approach. Given access to validation data for each target task, our method: (1) computes first-order gradient influence scores to measure how each training sample impacts task-specific performance, and (2) reaches influence consensus via majority voting to identify training samples that show consistent positive impact across multiple tasks. This consensus-based mechanism identifies universally valuable training examples: while some samples might be highly influential for individual tasks, we prioritize those that demonstrate broad utility across the task spectrum. While the computational cost of influence estimation is high, this front-loaded, one-time investment yields a standardized, compact dataset that can significantly accelerate development of multimodal models, and enables reusable gradient datastores that amortize costs across iterations and deliver long-term savings.
@@
Using ICONS, we create LLaVA-ICONS-133K, Cambrian-ICONS-1.4M, and Vision-Flan-ICONS-37K, automatically selected 20% subsets of LLaVA-665K [liu2024improved], Cambrian-7M [tong2024cambrian], and Vision-Flan-186K [visionflan] datasets, respectively. These compact datasets maintain 98.6%, 98.8%, and 99.8% of their original performance across multiple vision-language tasks, providing significant improvements over randomly selecting same-sized subsets (95.8%, 95.4%, and 91.6%) and eliminating approximately two-thirds of the performance drop from shrinking the training data. Moreover, our ICONS outperforms all baselines across different selection ratios, and remarkably achieves above-full-dataset performance, surpassing the original datasets at a 60% selection ratio for LLaVA-665K. Importantly, the selected subsets show strong transferability, e.g., LLaVA-ICONS-133K maintains 95.5-113.9% relative performance across unseen tasks, suggesting that ICONS identifies fundamentally valuable training data. We summarize our key contributions:
@@
    Our consensus-based selection outperforms all baselines (§4.2) and we ablate influence aggregation strategies and show the advantage of voting-based consensus (§4.3). ICONS exceeds 102% of full-dataset performance at a 60% selection ratio on LLaVA-665K4.5).
@@
    We release subsets LLaVA-ICONS-133K, Cambrian-ICONS-1.4M and Vision-Flan-ICONS-37K for resource-efficient development, demonstrating competitive performance including generalization to unseen tasks (§4.4) despite a 5x reduction in training data.
2 Related work
Visual instruction tuning.
Vision-language models (VLMs), e.g., Flamingo [flamingo], LLaVA [llava], BLIP2 [blip-2], and Cambrian [tong2024cambrian], enable understanding and reasoning across visual and textual modalities for various multimodal tasks. A key component in advancing VLMs is visual instruction tuning (VIT) [llava], a training process that enables these models to interpret and follow instructions within a vision-language context, transforming them into versatile multimodal assistants. This tuning process not only improves the models’ instruction-following abilities but also aligns their outputs more closely with user expectations, thus enhancing their utility in practical applications [llava]. To support this paradigm, recent VIT datasets have scaled to contain hundreds of thousands to millions of instruction-response pairs [liu2024improved, tong2024cambrian].
Data selection.
Data selection methods [survey] can be categorized based on the types of information they utilize for selection. Representation-based approaches [semdedup, coincide] leverage embeddings to capture data representations. Loss trajectory-based methods [mindermann2022prioritized] prioritize data points that contribute most significantly to reducing generalization error over training. Gradient-based techniques [el2n, xia2024less, deng2024influential] select data based on gradient information. Recent work has explored various approaches to select optimal visual instruction tuning datasets. Concurrent work TIVE [liu2024less] employs gradient-based selection to identify representative instances. TIVE assumes that the number of specialist data should be proportional to task difficulty and thus samples specialist data based on an estimation of task difficulty. Our method does not rely on this assumption – we directly select samples that benefit the greatest number of tasks. COINCIDE [coincide] clusters data based on representations associated with concept-skill compositions. Our work follows targeted instruction tuning selection approach LESS [xia2024less] to utilize gradient information to calculate the specialist influence (i.e., the influence on a specific task) and extends it to general scenarios by aggregating information across tasks and selecting data influential for multiple downstream tasks via majority voting.
3 ICONS: Influence CONsensus for vision-language data Selection
@@
Figure 2: ICONS. The Specialist stage (left) processes each task through: (1) warmup training on a small subset, (2) gradient computation for training and validation data, and (3) influence matrix computation for per-task scores. The Generalist stage (right) performs Influence Consensus to aggregate information across tasks. For each task, samples scoring above a percentile threshold (e.g., $80^{\text{th}}$ for 20% selection) receive a vote. Votes are summed across tasks, and samples with the highest vote counts are selected, creating a compact yet effective training dataset that performs well across all tasks.
@@ 3.1 Problem formulation
Given a large-scale visual instruction tuning dataset $\mathcal{D}=\{({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{N}$ containing $N$ samples, where each data point ${\bm{z}}_{i}=({\bm{I}}_{i},{\bm{x}}_{i},{\bm{y}}_{i})$ includes an image ${\bm{I}}_{i}$, natural language instruction ${\bm{x}}_{i}$, and corresponding target response ${\bm{y}}_{i}$ [^1] This framework supports multi-turn conversational data, yet we formalize the setup for single-turn instruction-tuning for clarity and simplicity., and given access to validation data $\mathcal{V}_{k}$ for each downstream task $T_{k}\in\mathcal{T}=\{T_{1},...,T_{K}\}$, our goal is to select a compact subset $\mathcal{S}\subset\mathcal{D}$ of size $M\ll N$ that maximizes model performance across multiple downstream tasks.
Concretely, we let $f_{\mathcal{S}}$ and $f_{\mathcal{D}}$ denote models trained on subset $\mathcal{S}$ and full dataset $\mathcal{D}$, respectively, and $\text{Score}(f,T_{k})$ be the task-specific evaluation score achieved by model $f$ on task $T_{k}$. $\text{Rel}(f_{\mathcal{S}},T_{k})=\frac{\text{Score}(f_{\mathcal{S}},T_{k})}{\text{Score}(f_{\mathcal{D}},T_{k})}$ quantifies the subset-trained model’s performance relative to that of the model trained on the entire dataset, with values close to 1 indicating that the subset maintains the performance of full training [coincide]. Our goal is then to identify the subset:
$\mathcal{S}^{*}=\operatorname*{arg\,max}_{\mathcal{S}\subset\mathcal{D},|\mathcal{S}|=M}\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k}).$ (1)
We define the average relative performance across all tasks as Rel. = $\sum_{k=1}^{K}\text{Rel}(f_{\mathcal{S}},T_{k})/K$, and our objective is to select a subset where Rel. $\approx 1$, i.e., the model trained on the subset achieves comparable performance to the one trained with full dataset.
@@ 3.2 Preliminaries
Building on the problem formulation in §3.1, we formalize how to estimate the influence of training samples on downstream task performance. Since our goal is to maximize $\text{Rel}(f_{\mathcal{S}},T_{k})$ across tasks as defined in Eqn. 1, we need an efficient way to estimate how each training sample contributes to the $\text{Score}(f_{\mathcal{S}},T_{k})$ term in the numerator. Denote a training data point as $\bm{z}$} and a validation data point as $\bm{z}$$^{\prime}$ from validation set $\mathcal{V}_{k}$ for task $T_{k}$. Following [pruthi2020estimating, xia2024less], we estimate how ${\bm{z}}\in\mathcal{D}$ affects validation loss by measuring its gradient alignment with reducing validation loss on $\mathcal{V}_{k}$, which directly impacts task-specific evaluation. When training with SGD and batch size 1, using data point $\bm{z}$} at timestep $t$ leads to a model update ${\bm{\theta}}_{t+1}={\bm{\theta}}_{t}-\eta_{t}\nabla\ell({\bm{z}};{\bm{\theta}}_{t})$, where $\eta_{t}$ is the learning rate. To reduce the computational cost, we use the first-order Taylor expansion to estimate the loss on a given validation data point $\bm{z}$$^{\prime}$ at time step $t+1$ as:
@@
Training sample $\bm{z}$}’s influence on validation sample $\bm{z}$$^{\prime}$ is:
$\mathcal{I}_{t}({\bm{z}}\rightarrow{\bm{z}}^{\prime})$ $=\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t+1})-\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t})$
$\approx-\eta_{t}\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle,$
which we refer to as an influence score. The gradient-based selection approach selects training samples $\{{\bm{z}}\}$ that maximize the gradient inner product $\langle\nabla\ell({\bm{z}}^{\prime};{\bm{\theta}}_{t}),\nabla\ell({\bm{z}};{\bm{\theta}}_{t})\rangle$ [^2] In practice, we use cosine similarity instead of direct inner products to avoid biasing selection toward shorter sequences, since gradient norms tend to be inversely correlated with sequence length as noted in [xia2024less]. through a greedy, first-order approximation, which leads to larger reductions in validation loss for point $\bm{z}$$^{\prime}$. While it omits second-order terms compared to influence functions [koh2017understanding], it provides an efficient approximation to rank the impact of training samples [xia2024less].
@@ 3.3 Specialist: individual task influence ranking
To rank the influence of training data for each target task, we compute the influence score of each training data point on a validation set that represents the target task distribution. Following LESS [xia2024less], the process involves three steps: (1) training the model on 5% randomly selected data as a lightweight warm-up to initialize visual instruction-following capabilities, (2) computing gradients for training and validation data and compressing the gradients via random projection, and (3) computing the influence score to quantify the impact of each training data on validation set.
Step 1: Warm-up Training.  Following LESS [xia2024less], we apply LoRA [hu2021lora] on a small random subset $\mathcal{D}_{\mathrm{warmup}}\subset\mathcal{D}$ (5%) to obtain $f_{\text{warmup}}=\text{LoRA}(f_{\text{base}},\mathcal{D}_{\text{warmup}})$, which develops basic visual instruction-following capabilities.
@@
where $y_{i}$ and $y^{\prime}_{j}$ are the targets for $z_{i}$ and $z^{\prime}_{j}$, respectively. In order to reduce computational and storage overhead, we apply random projection to the gradient feature: $\tilde{g}_{i}=Rg_{i}$ and $\tilde{g}^{\prime}_{j}=Rg^{\prime}_{j}$, where $R\in\mathbb{R}^{d^{\prime}\times d}$ is a random projection matrix with $d^{\prime}\ll d$ that preserves inner products with high probability [johnson1984extensions]. We further normalize the projected gradients, $\tilde{g}_{i}=\frac{\tilde{g}_{i}}{\|\tilde{g}_{i}\|_{2}},\tilde{g}^{\prime}_{j}=\frac{\tilde{g}^{\prime}_{j}}{\|\tilde{g}^{\prime}_{j}\|_{2}}$ to prevent bias from sequence length differences [xia2024less].
Step 3: Influence matrix computation.  We compute the influence matrix $I\in\mathbb{R}^{|\mathcal{D}|\times|\mathcal{D}_{\text{val}}^{k}|}$ where each entry $I_{ij}=\langle\tilde{g}_{i},\tilde{g}^{\prime}_{j}\rangle$ is the influence of the training data ${\bm{z}}_{i}$ on the validation data ${\bm{z}}^{\prime}_{j}$, following the setting in LESS [xia2024less], we take the average influence of training data ${\bm{z}}_{i}$ across all validation examples on the target task $k$ to obtain a representative task-level influence score:
@@
This influence estimation process provides a task-specific ranking for the training set $\mathcal{D}$ with respect to task $\mathcal{T}_{k}$, where a higher influence score $\bar{I}_{i}$ suggests a higher influence on $\mathcal{T}_{k}$. We can select a small training subset $\mathcal{S}_{k}$ for a given task $k$ by selecting the training examples ${\bm{z}}_{i}$ with the highest-influence values $\bar{I}_{k}({\bm{z}}_{i})$. This simple greedy approach has been shown by LESS to be successful, and thus we use it as our task-specific (“specialist”) baseline. However, recall that our goal is to select a single compact subset that maximizes the performance across all tasks. We address this disconnection between task-specific rankings and overall optimization by proposing a voting-based generalist approach to identify the most broadly impactful training data.